The side of a right triangle are 8,15 and 17 units. If each side is doubled, how many square units will be the area of the new triangle.
Answer by ankor@dixie-net.com(22740)

The side of a right triangle are 8,15 and 17 units.
If each side is doubled, how many square units will be the area of the new triangle.
:
We don't care about the hypotenuse (17)
:
Area of original triangle
.5 * 8 * 15 = 60
Area of side doubled triangle
.5 * 16 * 30 = 240
:
It will add: 240 - 60 = 180 sq units